OpenStudy (anonymous):

(e^2x)-(6e^x)+8=0 solve and give exact solutions

myininaya (myininaya):

\[\text{ let } u=e^x => u^2=e^{2x}\]

OpenStudy (anonymous):

let e^x be a a^2-6a+8=0

myininaya (myininaya):

yes that

OpenStudy (anonymous):

thanks